# Large Segmentation Design of Robust Photonic Quantum Gates

### v1

@@ III.1 Fidelity
and we focus on the range $1\%-10\%$, which is approximately around $10_{nm}-50_{nm}$. The fidelity loss term reads:
$\mathcal{L}_{F}=1-E_{\delta W\sim D}\left[\frac{1}{2}Tr\left(U_{ideal}^{\dagger}U\right)\right]\ ,$ (12)
@@ III.2 The Range of Width
where $\lambda_{W}$ is a hyperparameter for the minimal and maximal widths. We define the optimization domain $\Omega=(\Omega_{dz},\Omega_{w})$ as the set of all experimentally realizable waveguide widths. The number $N$ of segments in the design varies between 1 and 200, and we denote by $L$ the total length of the design. The degrees of freedom for the optimization are the set of widths and lengths of each segment: $\left\{W^{\text{L}}_{i},W^{\text{R}}_{i},dz_{i}\right\}_{1\leq i\leq N}$.
The waveguide width initialization is sampled uniformly in the range between the maximum and minimum of the feasible domain. Additionally, strong regularization is activated once the design is out of the optimization domain $\Omega$. It is also possible to add another regularization for the total design length $L$, but in our analysis we do not impose this restriction. Almost all the design lengths are between $10\mu m$ and $100\mu m$.
@@ III.3 Smooth Design Conditions
III.3.1 Baseline optimization
We define the baseline optimization as a search for the highest fidelity designs even with having abrupt changes between segments cross-section widths. This exhibits scattering and reflections, higher spatial modes generations and power loss, and requires a very high precision for manufacturing. Designs will be compared to the baseline ones.
III.3.2 Minimizing Fresnel Reflection
In order to reduce the power loss, we add a regularization term which reduces the back reflection coefficient. Thus, among the robust high-fidelity design parameter space, we search for the designs with the minimal power loss. The total cost function reads:
@@ III.3.2 Minimizing Fresnel Reflection
where $\mathcal{L}_{F},L_{\Omega},R$ are given by (12), (13) and (10), respectively.
III.3.3 Imposing An Adiabaticity Condition
@@
Figure 2: Example of two 100-segments designs of $\sqrt{X}$ gate in Silicon, with width noise standard deviation of $\sigma=30$nm ($\sigma_{\%}\approx 5\%$). Top: the Fresnel-based design with expectation fidelity of $0.91$ and reflection coefficient of ±$\pm 10^{-3}$. Bottom: the baseline with expectation fidelity of $0.9$ and reflection coefficient of ±$\pm 10^{-2}$. The roughness standard deviation (width change along the design) is $23$nm and $76$nm, respectively.
III.4 Error Correlations
The errors at different segments may be partially correlated. In order to quantify the effect on the design performance of different strength of correlations between the errors of different segments we introduce the error correlation matrix between the $k,l\in\{1,...N\}$ segments:
@@ III.4 Error Correlations
The dimensionless parameter $\mu$ controls the strength of the correlation: $\mu=0$ implies strong correlations between segments while $\mu\rightarrow\infty$ implies no correlation. The results of our correlations simulation will be discussed in the the simulation section (see IV.3).
III.5 Waveguide Material
We optimize the design for two types of materials: Silicon, which is widely used in photonic applications and has a relatively high refractive index (about 3.48), and LiNbO3 which is considered to be a promising candidate for photonic applications due to its outstanding optical properties, including a wide transparency window $(0.35\mu m-5\mu m)$, large optical nonlinearity and electro-optic coefficient, and strong acousto-optic effect. The Lithium Niobate has negative uniaxial birefringence, but we assume that only the first $TE_{01}$ mode is occupied.
For both materials we use a gap of $1.0~{}\mu m$ and the width of each waveguide is varied between $500~{}nm$ to $850~{}nm$. We further assume manufacture resolution (smallest resolution in the optimization) to be $20~{}nm$, waveguide height of $220~{}nm$, etch of $150~{}nm$ and wavelength of $1.55~{}\mu m$. To evaluate the mode-mismatch and coupling coefficients of the design we use Lumerical Eigensolver simulation (as shown in Fig. 1). We simulate different design geometries and build a look-up table to express the mode-mismatch coefficient $\Delta$ and the coupling coefficient $\kappa$ in terms of the geometrical waveguide parameters. The resolution of the geometry is $5~{}nm$, and we use a bi-linear interpolation to evaluate the values in-between.
@@ IV Optimization Results
In the optimization we divide each waveguide into $N=100$ segments. We use the PyTorch Package, with Adam optimizer, CosineAnnealing LR scheduler, with an initial learning rate of 0.1, batch size of 1024 samples, and the number of epochs was set to be around 5000. The device length is initialized randomly between $25~{}\mu m$ to $100~{}\mu m$. In our analysis we focus on the Pauli $X$ (without mode-mismatch) and the Hadamard $H$ (with mode-mismatch) single-qubit quantum gates. We found that other single-qubit gates show a similar behaviour. In Fig. 2 we see examples of composite designs (baseline and Fresnel based). Diagram of the trajectories on the Bloch sphere of designs for different optimization methods are shown in Fig. 3.
Figure 3: Bloch Sphere diagram of two optimization results. Left - Hadamard gate with 10 segments, initial state is $|0\rangle$ and final state is |+ket|$\rangle$| + ⟩. Right - Pauli X gate with 50 segments, initial state is $|+\rangle_{y}$ and final state is $|-\rangle_{y}$.
@@ IV.1 Mean Fidelity and Variance
Fig. 4 shows the expectation value and standard deviation of the fidelity for different numbers of segments taken in the range 1 to 200. To clean the optimization numerical noise, we avergated over 20 optimizations (for each number of segments). As can be clearly seen from the plot, increasing the number of segments leads to an oncrease in the gate fidelity, and the most significant improvement is obtained when the increasing the number of segments from 1 to 10. In Fig. 5 we plot the expectation value of the fidelity for different error magnitudes. Empirically, we find that the fidelity for Adiabatic and Fresnel designs are very similar to the baseline (see Fig. 9 and Appendix C for the list of the designs). We numerically observe large $N$ scaling laws for the expectation value and variance of the fidelity as functions of the number of segments:
@@
Figure 4: Mean Fidelity and standard deviation (shaded area) versus the number of segments for Pauli X, gate and Litium-Niobit material, assuming Normally distributed width noise with $\sigma=30$nm ($\sigma_{\%}\approx 5\%$). We see that the mean fidelity increases and variance decreases as the number of segments grows. This plot was created using an ensemble average over 20 optimizations for each $N$.
Figure 5: The expectation value of the fidelity for different number of segments, for increasing noise width. The design is optimized for $\sigma_{\%}\approx 5\%$. Increasing the number of segments leads to a higher robustness for large width errors.
@@ IV.2 Power Loss
To evaluate the power loss of the design, we use Lumerical simulator to measure the transmitted light for each transition between two neighbouring segments with different widths. In appendix B we present a comparison between different power loss calculation methods (Fresnel coefficients, FDTD, varFDTD and EME simulations), while here we focus on the comparison between the Fresnel coefficient (up to a global factor, see appendix B for details), and the FDTD simulation result (which is considered the most accurate among the simulation methods).
Fig. 6 shows the power loss for different number of segments, for the Hadamard gate made of Litium-Niobit material. While there are some differences between the Fresnel coefficients and the FDTD, their overall behaviour is very similar. We find empirically that optimizing over the Fresnel coefficients leads to a reduction in the total power loss estimated by the FDTD. To calculate the total power loss per design, we converted the loss per transition (between segments) using Eq. 8. Fig. 6 shows the averaged result over 10 optimizations.
Figure 6: Power loss versus number of segments for Hadamard, Litium-Niobit, using Fresnel coefficients (up to a global scale factor) in dashed line and FDTD simulation in solid line. We see an increase in the power loss as the number of segments increases, with an improvement when employing the Fresnel and the Adiabatic smoothing methods. This plot is created using an ensemble average over 10 optimizations for each $N$ and optimized for $\sigma_{\%}\approx 5\%$.
@@
Figure 7: Example of power density transmission cross-section curve using FDTD Lumerical simulator. The designs are of Silicon waveguides with $N=3$: Top - Baseline, Bottom - Fresnel-based. Green - Injected light in the beginning of the waveguide, blue - transmitted light to the end of the waveguide. The horizontal axis is the spatial cross-section of the waveguide, the vertical axis is the electromagnetic wave power density profile.
@@ IV.3 Fidelity and Error Correlation
When the correlation between the errors at different segments decreases, the composite pulse scheme for error mitigation becomes less effective YaronIdoNoiseThreshold . Here we analyze the impact of having a large number of segments when the error correlation decreases, where we model the error covariance matrix as (19). We plot in Fig. 8 the mean value and the standard deviation of the fidelity versus the error correlation strength $\mu$ (left, full correlation, and right, no correlation), for different numbers of segments for Pauli X gate, Silicon material. As expected, designs with a large number of segments show a better performance when the error correlations are small. We observe the same behaviour for other quantum gates.
@@
IV.4 Fidelity Versus Power Loss - An Effective Error Mitigation Region
Taking into account both the fidelity of the quantum gate and the power loss of the design, we plot in Fig. 9 the mean fidelity versus the power loss and identify the effective error mitigation region. The designs are listed in table 1 in appendix C
@@ V Discussion and Outlook
We studied the performance of the segmented composite pulse error mitigation scheme applied to the design of photonic quantum gates, when increasing the number of segments $N$ and varying the error correlation strength $\mu$. We considered the fidelity and the power loss of single-qubit quantum gates realized by optical waveguides in the dual rail framework. We quantified the impact of increasing $N$ on the fidelity and the power loss of the quantum gates and observed certain asymptotic scaling laws. Empirically we found that $O(10)$ number of segments yields effective designs.
To reduce the power loss when increasing $N$, we applied two optimization regularizations by imposing constraints on the Fresnel reflection coefficient and an adiabatic condition on the waveguides width profile. We have not identified a significant difference in the gate performance between the two regularizations. However, from practical considerations it is simpler to apply the adiabatic one since it does not require a pre-calculation of the effective refraction index in the waveguide.
@@
We studied the effect of varying the correlation strength between the errors of different segments. Large correlations between the errors imply better gate fidelity, since the different segments can effectively compensate the accumulated errors. We showed that a large number of segments allows to overcome the problem of weak error correlations.
There are several directions to follow. One is to analyze the continuum limit of the large $N$ designs that result in smooth designs. A second one is to validate our error model against experimental data. A third direction to follow is to study the robust designs for QIP systems that include several or many quantum gates.
@@
Acknowledgements We would like to thank J. Drori and Y. Piasetzky for their assistance in extracting parameters using Lumerical simulations. This work is supported in part by the Israeli Science Foundation Excellence Center, the US-Israel Binational Science Foundation, and the Israel Ministry of Science.
Appendix A Single Qubit Gates
The time evolution of a two-level systems (qubit) is governed by the Schrödinger equation:
$i\partial_{t}\left(\begin{array}[]{l}c_{1}(t)\\ c_{2}(t)\end{array}\right)=\left(\begin{array}[]{cc}-\Delta(t)&\kappa^{*}(t)\\ \kappa(t)&\Delta(t)\end{array}\right)\left(\begin{array}[]{l}c_{1}(t)\\ c_{2}(t)\end{array}\right)\ ,$ (28)
where $c_{1}(t)$ and $c_{2}(t)$ are the probability amplitudes at time $t$ of being in the states $|0\rangle$ and $|1\rangle$, respectively, $\kappa(t)$ is the (complex) Rabi frequency, $\Delta(t)$ is the (real) mode-mismatch, and we set $\hbar=1$. The unitary propagator of such a system is:
$U(t,0)=\mathcal{T}\left[\exp\left[-i\int_{0}^{t}\left(\begin{array}[]{cc}-\Delta\left(t^{\prime}\right)&\kappa^{*}\left(t^{\prime}\right)\\ \kappa\left(t^{\prime}\right)&\Delta\left(t^{\prime}\right)\end{array}\right)\right]dt^{\prime}\right]\ .$ (31)
When $\kappa$ and $\Delta$ are time independent, the propagator simplifies to
$U(t,0)=\exp\left[-it\left(\begin{array}[]{cc}-\Delta&\kappa^{*}\\ \kappa&\Delta\end{array}\right)\right]\ .$ (34)
Physical systems that exhibit this $SU(2)$ dynamics implement one-qubit gates as well as probabilistic two-qubit gates. In silicon photonics, the analog of the time is the propagation axis, and the coupling coefficient $\kappa$ is real. Therefore, for a single segment of length $z$:
$U(\kappa,\Delta,z)=\exp\left[-iz\left(\kappa X-\Delta Z\right)\right]\ ,$ (35)
where $X$ and $Z$ are the Pauli matrices. We can rewrite the $SU(2)$ quantum gate as:
$U(\kappa,\Delta,z)=\exp\left[-iz\Omega_{g}\left(\frac{\kappa}{\Omega_{g}},0,\frac{\Delta}{\Omega_{g}}\right)\cdot\vec{\sigma}\right]$ (36)
$=\cos\left(z\Omega_{g}\right)I-\frac{\kappa}{\Omega_{g}}\sin\left(z\Omega_{g}\right)X-\frac{\Delta}{\Omega_{g}}\sin\left(z\Omega_{g}\right)Z\ ,$ (37)
where $\Omega_{g}\equiv\sqrt{\kappa^{2}+\Delta^{2}}$.
@@ Appendix B Power Loss Simulation Comparison
In Fig. 10 we show a comparison between the power loss simulator (left) to the Fresnel coefficient (right). We consider the FDTD the most accurate simulation method and this is the reason we present this method along with the Fresnel coefficients. The Fresnel coefficient approximation of the power loss compared to the simulation, are related by a global factor of about 5 (see 6 in the main text).
Figure 10: Our power loss simulation using EME (top-left), Fresnel coefficients (top-right), varFDTD (bottom-left) and FDTD (bottom-right). The simulation evaluated the power loss due to a single abrupt transition from width $W_{1}$ (x axis in nm) to width $W_{2}$ (y axis in nm), for Silicon and $H=220nm$.

### v2

@@ III.1 The Fidelity Term
and we focus on the range $1\%-10\%$, which is approximately $10_{nm}-50_{nm}$. The fidelity loss term reads:
$\mathcal{L}_{F}=1-E_{\delta W\sim D}\Big{[}F\left(U_{\text{ideal}},U\right)\Big{]}\ ,$ (24)
@@ III.2 The Range of Width and Length
where $\lambda_{W}$ is a hyperparameter for the minimal and maximal widths. We define the optimization domain $\Omega=(\Omega_{dz},\Omega_{w})$, as the set of all experimentally realizable waveguide widths. The number $N$ of segments in the design varies between 1 and 200, and we denote by $L$ the total length of the design. The degrees of freedom for the optimization are the set of widths and lengths of each segment: $\left\{W^{\text{L}}_{i},W^{\text{R}}_{i},dz_{i}\right\}_{1\leq i\leq N}$.
The waveguide width initialization is sampled uniformly in the range between the maximum and minimum of the feasible domain. Furthermore, strong regularization is activated once the design is out of the optimization domain $\Omega$.
To avoid overly short segments that could violate our assumptions and necessitate high fabrication resolution, we restricted our focus to devices with sufficiently large segment lengths. Specifically, we ensured that the dimensionless parameter characterizing propagation, defined as the ratio of each segment’s length to the wavelength in the material, remains sufficiently large.
While it is also possible to impose an additional regularization on the total design length $L$, this restriction is not applied in our analysis. Instead, filtered-out devices with a minimum segment length less than $1\mu m$. Our results show that nearly all designs naturally exhibit total lengths in the range of $10\mu m$ to $100\mu m$ (see the table of our designs 1 in the Appendix).
@@ III.3 Smooth Design Conditions
We define the baseline optimization, as a search for the highest fidelity designs, without taking into account any smoothness requirement. This lead to designs with high back-scattering, and power loss. The alternative optimized designs will be evaluated in comparison to these baseline configurations.
III.3.1 Minimizing Fresnel Reflection
In order to reduce the power loss, we add a regularization term, which reduces the back reflection coefficient. Thus, among the robust high-fidelity design parameter space, we search for the designs with the minimal power loss. The total cost function reads:
@@ III.3.1 Minimizing Fresnel Reflection
where $\mathcal{L}_{F},L_{\Omega},R$ are given by Eq. 24, 25 and 21, respectively.
III.3.2 Imposing An Adiabaticity Condition
@@
III.4 Segments Error Correlation Model
We also account for cases with partial correlation, where errors in different segments may exhibit varying degrees of dependency. To assess the impact of correlation strength on design performance, we introduce an error correlation matrix to model the relationships between errors across $k,l\in\{1,...N\}$ segments:
@@ III.4 Segments Error Correlation Model
The dimensionless parameter $\mu$ controls the strength of the correlation: $\mu=0$ implies strong correlations between segments while $\mu\rightarrow\infty$ implies no correlation. The results of our correlations simulation will be discussed in the simulation section (see sub-section IV.3).
We optimize the design for two materials: Silicon and $\text{LiNbO}_{3}$ (Lithium Niobate). Silicon, widely used in photonic applications, is notable for its high refractive index of approximately 3.48. $\text{LiNbO}_{3}$, on the other hand, is a promising candidate due to its exceptional optical properties, including a wide transparency window ($0.35\mu m$ to $5\mu m$), large optical nonlinearity, high electro-optic coefficient, and strong acousto-optic effect. We assume in our analysis that only the fundamental $TE_{01}$ mode is occupied.
For both materials we use a gap of $1.0~{}\mu m$ and the width of each waveguide is varied between $500~{}nm$ to $850~{}nm$. We further assume the manufacture resolution (smallest resolution in the optimization) to be $20~{}nm$, waveguide height of $220~{}nm$, etch of $150~{}nm$ and wavelength of $1.55~{}\mu m$. To evaluate the mode-mismatch and coupling coefficients of the design we use Lumerical Eigensolver simulation (as shown in Fig. 2).
@@ IV Optimization Results
In the optimization we divide each waveguide into $N=100$ segments. We use the PyTorch Package, with Adam optimizer, CosineAnnealing LR scheduler, with an initial learning rate of 0.1, batch size of 1024 samples, and the number of epochs was set to be around 5000. The length of the device is randomly initialized between $25~{}\mu m$ and $100~{}\mu m$. In our analysis we focus on the Pauli $X$ (without mode-mismatch) and the Hadamard $H$ (with mode-mismatch) single-qubit quantum gates.
In this study, we observed that the scaling behavior of the two materials is quite similar, as well as the gate types. Consequently, we decided to concentrate on presenting our analysis for only one material, and one gate at a time.
To visually represent the geometric design achieved through the optimization process, refer to Fig. 6. For a depiction of the trajectories on the Bloch sphere, please see Appendix A.
@@ IV.1 Mean Fidelity and Variance
Fig. 3 shows the expectation value and standard deviation of the fidelity for different numbers of segments taken in the range 1 to 200. The presented value for each segment is an average over 20 optimizations values. Increasing the number of segments leads to an increase in the gate fidelity. The most significant improvement is obtained when the increasing the number of segments from 1 to 10. In Fig. 4 we show the expectation value of the fidelity for different error magnitudes. Empirically, we find that the fidelity for Adiabatic and Fresnel designs are very similar to the baseline. We numerically observe large $N$ scaling laws for the expectation value and variance of the fidelity as functions of the number of segments:
@@
Figure 3: Mean Fidelity and standard deviation (shaded area) versus the number of segments for $\text{LiNbO}_{3}$ X gate, assuming Normally distributed width noise with $\sigma=30_{nm}$ ($\sigma_{\%}\approx 5\%$). We see that the mean fidelity increases and variance decreases as the number of segments grows. This plot was created using an ensemble average over 20 optimizations for each $N$.
Figure 4: The expectation value of the fidelity for different number of segments, for increasing noise width, in $\text{LiNbO}_{3}$ X gate. The design is optimized for $\sigma_{\%}\approx 5\%$. Increasing the number of segments leads to a higher robustness for large width errors.
@@ IV.2 Power Loss
To evaluate the power loss of the design, we use Lumerical simulator to measure the transmitted light for each transition between two neighboring segments with different widths. In appendix B we present a comparison between different power loss calculation methods (Fresnel coefficients, FDTD, varFDTD and EME simulations), while here we focus on the comparison between the Fresnel coefficient (up to a global factor, see appendix B for details), and the FDTD simulation result, which is considered the most accurate among the simulation methods.
Fig. 5 shows the power loss for different number of segments, for the Hadamard gate made of $\text{LiNbO}_{3}$, each data point is averaged over 10 optimizations.. While there are some differences between the Fresnel coefficients and the FDTD, their overall behavior is very similar. We find empirically that optimizing over the Fresnel coefficients leads to a reduction in the total power loss estimated by the FDTD. To calculate the total power loss per design, we converted the loss per transition (between segments) using Eq. 19.
Figure 5: Power loss versus number of segments for Hadamard, in $\text{LiNbO}_{3}$ based Hadamard-Gate, using Fresnel coefficients (up to a global scale factor) in dashed line and FDTD simulation in solid line. We see an increase in the power loss as the number of segments increases, with an improvement when employing the Fresnel and the Adiabatic smoothing methods. This plot is created using an ensemble average over 10 optimizations for each $N$ and optimized for $\sigma_{\%}\approx 5\%$.
@@
Figure 6: An illustration comparing baseline (top) and Fresnel-based (bottom) optimizations. On the left, the power density transmission cross-section is depicted using the FDTD Lumerical simulator, while on the right, the geometrical configuration along the propagation axis is shown. The green line corresponds to the intensity profile at the waveguide’s entry, whereas the blue line indicates the final intensity profile. This configuration consists of 100 segments, using Silicon and a $\sqrt{X}$ gate, with a width noise standard deviation of $\sigma=30_{nm}$ ($\sigma_{\%}\approx 5\%$). Both designs achieve a fidelity of $0.91$.
@@ IV.3 Fidelity and Error Correlation
When the correlation between the errors at different segments decreases, the composite pulse scheme for error mitigation becomes less effective YaronIdoNoiseThreshold . Here we analyze the impact of having a large number of segments when the error correlation decreases, where we model the error covariance matrix as Eq. 31. We plot in Fig. 7 the mean value and the standard deviation of the fidelity versus the error correlation strength $\mu$ (left, full correlation, and right, no correlation), for different numbers of segments for the Pauli X gate, Silicon material. As expected, designs with a large number of segments show better performance when the error correlations are small. We observe the same behavior for other quantum gates.
@@
IV.4 Fidelity Versus Power Loss, and An Effective Error Mitigation Region
Taking into account both the fidelity of the quantum gate and the power loss of the design, we plot in Fig. 8 the mean fidelity versus the power loss. As seen, fidelity improves with a higher number of segments. While power loss also rises with more segments, it can be mitigated by incorporating the proposed smooth-design regularization. The dashed line outlines a region we term as the Effective Error Mitigation Region, as it encompasses exclusively designs that achieve both high fidelity and minimal power loss. All the designs presented in this figure are listed in table 1 in appendix C.
@@ V Summary and Outlook
We studied the performance of the segmented composite pulse error mitigation scheme applied to the design of photonic quantum gates, when increasing the number of segments $N$ and varying the error correlation strength $\mu$. We quantified the impact of increasing $N$ on the fidelity and the power loss of the quantum gates and observed certain asymptotic scaling laws. Empirically we found that $O(10)$ number of segments yields effective designs.
To minimize power loss with increasing $N$, we applied two optimization regularization: one imposing constraints on the Fresnel reflection coefficient and the other introducing an adiabatic condition on the waveguide width profile. Both approaches yielded comparable gate performance, although the adiabatic condition proves to be slightly more convenient within the optimization context.
@@
We also examined the impact of varying correlation strength between errors in different segments. Our results showed that stronger correlations improve gate fidelity, as the segments can better compensate for accumulated errors. Notably, a large number of segments mitigates the challenges associated with weak error correlations.
There are several natural research directions to follow. First, while our results are specifically tailored to dual-rail photonic gates, the insights gained from this study are expected to generalize across various noise models and error frameworks, including those with different error correlations. These findings hold potential for application to other quantum systems and platforms, providing valuable guidance for error mitigation strategies in broader quantum contexts.
Another important direction to follow is to study robust designs for QIP systems that include several or many quantum gates. In such cases, the optimization framework is likely to find robust designs that are better than a simple combination of individually robust single-qubit gates.
@@
Acknowledgements We would like to thank J. Drori and Y. Piasetzky for their assistance in extracting parameters using Lumerical simulations. This work is supported in part by the Israeli Science Foundation Excellence Center No. 2312/21, the US-Israel Binational Science Foundation, and the Israel Ministry of Science.
Appendix A Bloch Sphere Trajectories
Fig. 9 shows a trajectory comparison for two different cases.
Figure 9: Bloch Sphere diagram of two configurations in Silicon. Left - Hadamard gate with 10 segments, initial state is $|0\rangle$ and final state is |+ket|$\rangle$| + ⟩. Right - Pauli X gate with 50 segments, initial state is $|+\rangle_{y}$ and final state is $|-\rangle_{y}$.
@@ Appendix B Power Loss Simulation Comparison
In Fig. 10 we show a comparison between the power loss simulator (left) to the Fresnel coefficient (right). We consider the FDTD the most accurate simulation method and this is the reason we present this method along with the Fresnel coefficients. The Fresnel coefficient approximation of the power loss compared to the simulation, are related by a global factor of about 5 (see Fig. 5 in the main text).
Figure 10: Our power loss simulation using EME (top-left), Fresnel coefficients (top-right), varFDTD (bottom-left) and FDTD (bottom-right). The simulation evaluated the power loss due to a single abrupt transition from width $W_{0}$ (x axis in nm) to width $W_{1}$ (y axis in nm), for Silicon and $H=220nm$.